For a fixed fee of $30, she is allowed 100 minutes of free calls. For any calls above 100 minutes, she pays an additional fixed fee of $20 as well as a per-minute charge for each additional minute above 100 minutes. the per-minute charge is 20 cents between 100 and 600 minutes, and 60 cents after 600 minutes.

1. Let B be the monthly bill in dollars and T be the total time of calls in minutes. Write an equation that describes Kayla's monthly bill from T = 0 to T = 1000 minutes

2. Construct a graph that describes Kayla's monthly bill as a function of total time.
Answer #1

(1) B = 30 for t ≤ 100,  = 50 + 0.20t for 100 < t ≤ 600,  = 50 + 0.60t for 600 < t ≤ 1000

(2) Graph:

700 650 600 550 500 450 400 350 300 250 200 150 100 50 0 100 200 300 400 500 600 700 800 900 1000 Minutes